# Quest: Query-centric Data Synthesis Approach for Long-context Scaling of Large Language Model

### v3

@@ 1 Introduction
Figure 1: (Left) Distribution of long documents up to 128k in Pile. (Right) Distribution of 128k long-context data synthesized by Quest.
@@
To obtain long-context data, Xiong et al. (2023); Fu et al. (2024) have filtered out long documents meeting the target context length, though these documents often derive from a few specific domains. As shown in Figure 1 (left), our analysis of the widely used Pile corpus Gao et al. (2020) reveals that long documents are mainly concentrated in the Books3 dataset, leading to a skewed distribution that worsens with longer target contexts. Previous studies Guu et al. (2020); Levine et al. (2021); Shi et al. (2023) have suggested synthesizing long-context data by aggregating semantically similar documents, such as concatenating a document with its top k retrieved documents. However, these methods often result in redundancy due to similar sentences, especially in large-scale corpora, reducing token prediction difficulty and context diversity, thereby weakening long-context modeling effectiveness (see Section 6.2 for analysis). Thus, there’s a pressing need for a method to effectively aggregate relevant but low-redundant documents for long-context data synthesis. Additionally, this method must be highly scalable to construct large datasets for continued training.
@@ 2 Related Work
The success of LLMs has sparked interest in enabling these models to process longer texts. Some works adapt models for longer texts without additional training by modifying position encoding. For example, Han Han et al. (2023) and Xiao Xiao et al. (2023) adjust the attention matrix to generate long contexts, while Jin Jin et al. (2024) compresses position encoding into the pre-trained range. Other works involve continued training for better performance. Xiong Xiong et al. (2023) demonstrates that long-context capabilities can be acquired by continually pre-training from short-context models. Chen Chen et al. (2023b) uses position interpolation to change the distribution of position encoding, and Yen Yen et al. (2024) proposes context expansion with parallel encoding. Advances have also been made using RoPE Su et al. (2021), enabling LLMs to handle longer positions. PoSE Zhu et al. (2023) employs skip-wise position indices, allowing position encoding to adapt to different lengths. However, these approaches often overlook the scarcity and uneven distribution of long text data during continued training, relying on filtering long documents from existing corpora Xiong et al. (2023); Fu et al. (2024) or randomly splicing short documents to achieve a fixed length Roziere et al. (2023); Chen et al. (2023c); Tworkowski et al. (2024); Chen et al. (2023a); Li et al. (2023).
@@
Acquiring effective long-context data for training is challenging. Some previous retrieval-augmented pre-training works Guu et al. (2020); Levine et al. (2021) can synthesize long-context data. Guu Guu et al. (2020) clusters semantically similar texts within the same context window, while Levine Levine et al. (2021) shows that incorporating semantically related but non-adjacent sentences within the same pre-training example enhances sentence representations. Shi Shi et al. (2023) uses a traveling salesman algorithm to address document redundancy in the kNN method. However, previous data synthesis efforts were limited to context lengths of 8k or less, and the benefits of synthesizing longer data were unclear. This work uses various methods to synthesize texts up to 128k, demonstrating the effectiveness of Quest in synthesizing long-context data.
@@ 3 Method
This section details our proposed Query-centric Data Synthesis (Quest) method. Given a dataset with diverse documents $D=\{d_{i}\}$, our goal is to effectively aggregate relevant but low-redundant documents for synthesizing training texts with a context length of $L$. An overview of our approach can refer to Figure 2. Quest mainly includes five steps. First, a query $\{q_{i}\}$ is predicted for each document $\{d_{i}\}$ in the corpus. Next, a topic keyword $\{k_{i}\}$ is extracted from each query. Thirdly, documents with the same keyword are grouped or indexed together. Then, we split the keyword-based inverted indexes according to the number of documents. During training, at each step, we perform sampling without replacement for the documents $\{k_{i}\}$ within a sampled index. The details of each stage are provided below.
@@
    Query Prediction: We utilize the open-source doc2query modelNogueira et al. (2019) to predict queries $\{q_{i}\}$ for each document $\{d_{i}\}$. For texts that exceed the context length limit of the doc2query model, we segment them into parts and generate a query for each segment. Consequently, for a document $\{d_{i}\}$, a list of queries $Q_{i}=\{q_{i}^{1},\dots,q_{i}^{n}\}$ is predicted.
@@
    Indexes Split: We rank the keyword-based inverted indexes in ascending order based on the number of documents within each index and divide the sorted indexes into two sets. The top-ranked split_ratioof the keyword-based inverted indexes are assigned to the short-index set $I_{s}$, while the remainder is assigned to the long-index set $I_{l}$.
@@ 4 Experiments
4.1 Experimental Setup
We apply the Quest method on Pythia’s pre-training data, i.e., the Pile dataset, which does not lead to domain transfer issues. Specifically, we extract 30B tokens of keyword-indexed documents from the 300B tokens of the original Pile dataset. Documents indexed by an identical keyword are randomly concatenated to form long-context data that reach the training context length.
During training, we use the open-source framework GPT-NeoX [^2] https://github.com/EleutherAI/gpt-neox with a batch size of 4M tokens for all settings. The AdamW optimizerLoshchilov and Hutter (2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ parameters and a cosine learning rate schedule is employed. We also use Flash Attention2Dao (2023) and ZeRORajbhandari et al. (2020) to optimize memory and performance. The learning rates are $5e^{-5}$ for the 1.4B model, $4e^{-5}$ for the 6.9B model, and $2e^{-5}$ for the 12B model. For more details, see Appendix B.1.
4.2 Baselines Methods
  1. 1.
  2. 2.
  3. 3.
    ICMLShi et al. (2023) Method is a recently proposed method that utilizes a traveling salesman algorithm to alleviate the document redundancy problem in the kNN method by ranking similarities and determining the optimal training path.
@@ 4.3 Evaluation
    Long-text Benchmark: For 32k context length, we adopt the widely-used LongbenchBai et al. (2023) Benchmark, testing six task types: Single-document QA, Multi-document QA, Summarization, Few-shot learning, Synthetic, and Code completion, totaling 17 datasets. For 128k context length, we focus on the widely-used Longbook QA taskZhang et al. (2024), on which the pre-trained models perform reasonably well without instruction tuningFu et al. (2024).
@@
Table 1 presents the comparison results on the Longbench, showing that Quest outperforms other methods across various model sizes. The results indicate that KNN and ICLM underperform the Standard method, likely due to grouping textually similar documents, which leads to redundancy in the context. To further analyze the behavior of these methods with a 32k context length, we use TSNE for visualization. Figure 3 (left) demonstrates that the Standard method results in the most dispersed document clusters due to random aggregation. In contrast, KNN and ICLM methods show tightly clustered documents, while Quest exhibits moderate aggregation. This suggests that Quest minimizes document redundancy within the extended context, aligning with its superior performance in the Longbench results.
@@
Figure 3: TSNE visualization of aggregated documents from different methods. See Appendix D for more examples.
Quest retains good performance on short text.
@@
Figure 4: Performance on Needle-in-a-Haystack task. The x-axis depicts the document’s length, termed the "haystack," whereas the y-axis illustrates the position of the "needle" (a brief sentence) within the document, spanning from 1K to 128K tokens.
4.4 Performance on the SOTA Model.
Figure 5: Unlike Figure 4, in the scenario of a 500k long text, the "needle" is a string of random numbers.
@@ 5 Scaling Law of Synthesized Long-context Data
Furthermore, we verify the correctness of the learned scaling law formula under the data size of 8 billion. For each model size, we compare the relative error between the validation loss at the end of training with 8 billion data and its predicted counterpart via the learned scaling law. The relative error of the 1.4B model is 0.5%, the relative error of the 6.9B model is -0.5%, and the relative error of the 12B model is 0.4%. These negligible errors are a strong demonstration of the scalability and predictability of Quest’s data synthesis approach.
6 Analysis
This section provides an in-depth analysis of the Quest method. Considering the high computational cost of LLM experiments, our ablation experiments are performed under a 32k context length, 1.4B model size setting unless otherwise stated.
6.1 Quest’s Advantage Gradually Expands with Training Progress
@@ 6.3 Quest-Synthesized Long-context Outperforms Long Document
6.4 Impact of Split Ratio
This section studies the impact of split_ratio, which controls the proportion of oversampled keyword-based inverted indexes. Figure 9 shows the performance changing trend with split_ratio increase. Experimental results indicate that overall performance follows a trend of initially growing and then declining with the increase in split_ratio. The proportion of oversampled indexes between 10% and 30% yields the best results in average. This shows that appropriate oversampling of the indexes with fewer documents is beneficial to long-context modeling, again illustrating the importance of balanced data distribution for long-context capabilities.
@@ B.2 Filtering of Keywords
Figure 12: Visualizing Documents Comprising a 128k Context.

### v4

@@ 1 Introduction
Figure 1: (Left) Distribution of long documents exceeding 128k tokens in the Pile dataset. (Right) Distribution of 128k token long-context data synthesized by Quest. By employing Quest, we successfully increased the domain diversity of the long-context training set.
@@
To obtain long-context data, Xiong et al. (2023) and Fu et al. (2024) have filtered out long documents meeting the target context length, though these documents often derive from a few specific domains. As shown in Figure 1 (left), our analysis of the widely used Pile corpus Gao et al. (2020) reveals that long documents are mainly concentrated in the Books3 dataset, leading to a skewed distribution that worsens with longer target contexts. Previous studies Guu et al. (2020); Levine et al. (2021); Shi et al. (2023) have suggested synthesizing long-context data by aggregating semantically similar documents, such as concatenating a document with its top k retrieved documents. However, these methods often result in redundancy due to similar sentences, especially in large-scale corpora, reducing token prediction difficulty and weakening long-context modeling effectiveness (see Section 6.2 for analysis). Thus, there’s a pressing need for a method to effectively aggregate relevant but low-redundant documents for long-context data synthesis. Additionally, this method must be highly scalable to construct large datasets for continued training.
@@ 2 Related Work
The success of LLMs has sparked interest in enabling these models to process longer texts. Some works adapt models for longer texts without additional training by modifying position encoding. For example, Han et al. (2023) and Xiao et al. (2023) adjust the attention matrix to generate long contexts, while Jin et al. (2024) compresses position encoding into the pre-trained range. Other work involves continued training for better performance. Xiong et al. (2023) demonstrates that long-context capabilities can be acquired by continually pre-training from short-context models. Chen et al. (2023b) uses position interpolation to change the distribution of position encoding, and Yen et al. (2024) proposes context expansion with parallel encoding. Advances have also been made using RoPE Su et al. (2021), enabling LLMs to handle longer positions. PoSE Zhu et al. (2023) employs skip-wise position indices, allowing position encoding to adapt to different lengths. However, these approaches often overlook the scarcity and uneven distribution of long text data during continued training, relying on filtering long documents from existing corpora Xiong et al. (2023); Fu et al. (2024) or randomly splicing short documents to achieve a fixed length Roziere et al. (2023); Chen et al. (2023c); Tworkowski et al. (2024); Chen et al. (2023a); Li et al. (2023).
@@
Acquiring effective long-context data for training is challenging. Some previous retrieval-augmented pre-training works Guu et al. (2020); Levine et al. (2021) can synthesize long-context data. Guu et al. (2020) clusters semantically similar texts within the same context window, while Levine et al. (2021) shows that incorporating semantically related but non-adjacent sentences within the same pre-training example enhances sentence representations. Shi et al. (2023) uses a traveling salesman algorithm to address document redundancy in the kNN method. However, previous data synthesis efforts were limited to context lengths of 8k or less, and the benefits of synthesizing longer data were unclear. This work uses various methods to synthesize texts up to 128k, demonstrating the effectiveness of Quest in synthesizing long-context data.
@@ 3 Method
This section details our proposed Query-centric Data Synthesis (Quest) method. Algorithm 1 presents the method for synthesizing long-context data. Given a dataset with diverse documents $D=\{d_{i}\}$, our goal is to effectively aggregate relevant but low-redundant documents for synthesizing training texts with a context length of $L$. An overview of our approach can refer to Figure 2. Quest mainly includes five steps.
@@
    Query Prediction: We utilize the open-source doc2query model Nogueira et al. (2019) to predict queries $\{q_{i}\}$ for each document $\{d_{i}\}$. We segment texts that exceed the context length limit of the doc2query model into parts and generate a query for each segment. Consequently, for a document $\{d_{i}\}$, a list of queries $Q_{i}=\{q_{i}^{1},\dots,q_{i}^{n}\}$ is predicted.
@@
    Indexes Split: We found that the number of documents associated with different keywords varies significantly. To address this imbalance and achieve a more balanced data distribution, we implement oversampling for documents with less frequent keywords. Specifically, we rank the keyword-based inverted indexes in ascending order based on the number of documents within each index and divide the sorted indexes into two sets. The top-ranked split_ratioof the keyword-based inverted indexes are assigned to the short-index set $I_{s}$, while the remainder is assigned to the long-index set $I_{l}$.
@@ 4 Experiments
4.1 Experimental Setup
We apply the Quest method on Pythia’s pre-training data, i.e., the Pile dataset, which does not lead to domain transfer issues. Specifically, we extract 30B tokens of keyword-indexed documents from the 300B tokens of the original Pile dataset. Documents indexed by an identical keyword are randomly concatenated to form long-context data that reach the training context length. We use identical training sets for all methods to ensure a strictly fair comparison. The only variation between the methods is rearranging documents into long-context data.
During training, we use the open-source framework GPT-NeoX [^2] https://github.com/EleutherAI/gpt-neox with a batch size of 4M tokens for all settings. The AdamW optimizer Loshchilov and Hutter (2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ parameters and a cosine learning rate schedule is employed. We also use Flash Attention2 Dao (2023) and ZeRO Rajbhandari et al. (2020) to optimize memory and performance. The learning rates are $5e^{-5}$ for the 1.4B model, $4e^{-5}$ for the 6.9B model, and $2e^{-5}$ for the 12B model. For more details, see Appendix B.
4.2 Baselines Methods
  1. 1.
  2. 2.
  3. 3.
    ICLM Shi et al. (2023) Method is a recently proposed method that utilizes a traveling salesman algorithm to alleviate the document redundancy problem in the kNN method by ranking similarities and determining the optimal training path.
@@ 4.3 Evaluation
    Long-text Benchmark: For 32k context length, we adopt the widely-used Longbench Bai et al. (2023) Benchmark, testing six task types: Single-document QA, Multi-document QA, Summarization, Few-shot learning, Synthetic, and Code completion, totaling 17 datasets. For 128k context length, we focus on the widely-used Longbook QA task Zhang et al. (2024b), on which the pre-trained models perform reasonably well without instruction tuning Fu et al. (2024).
@@
Table 1 compares the Longbench results, showing that Quest consistently outperforms other methods across model sizes, especially on the Code dataset. We attribute this to the use of Levenshtein distance, a more stringent and discriminative metric than F1 score or ROUGE, as it accounts for character sequence.
The results also show that KNN and ICLM underperform compared to the Standard method, likely due to redundancy from grouping similar documents. To analyze this further with a 32k context length, we used TSNE visualization. Figure 3 (left) shows that the Standard method yields the most dispersed clusters due to random aggregation, while KNN and ICLM produce tightly clustered documents. In contrast, Quest shows moderate clustering, indicating it reduces document redundancy, which aligns with its superior Longbench performance.
Figure 3: TSNE visualization of aggregated documents from different methods. Quest maintains balanced document distribution across varying context lengths. See Appendix D for more examples.
@@
Quest retains good performance on short text.
@@
4.4 Performance on the SOTA Model.
Figure 4: Performance on the Needle-in-a-Haystack task. The x-axis shows the document length ("haystack"), while the y-axis indicates the position of the "needle" (a short sentence) within the document, ranging from 1K to 128K tokens. "Acc" represents the percentage of model responses rated as fully accurate (scoring 10 in GPT-4’s evaluation) out of the total responses generated.
Figure 5: To facilitate the testing of ultra-long texts, we followed the approach used in LongVA Zhang et al. (2024a) and LWM Liu et al. (2024) for the Needle-in-a-Haystack task, where the goal is to recall a random numeric string.
@@ 5 Scaling Law of Synthesized Long-context Data
We validated the learned scaling law on an 8 billion data size by comparing the relative error between each model’s final validation loss and its predicted value. The relative errors were 0.5% for the 1.4B model, -0.5% for the 6.9B model, and 0.4% for the 12B model, demonstrating the scalability and accuracy of Quest’s data synthesis approach with minimal deviation.
@@
6 Analysis
We tested 26 datasets of varying lengths and types, with Quest consistently ranking in the top 1 or 2, making it the overall best method. This section provides an in-depth analysis of Quest. Due to the high computational cost of LLM experiments, our ablation studies are conducted with a 32k context length and 1.4B model size, unless otherwise noted.
6.1 Quest’s Advantage Gradually Expands with Training Progress
@@ B.1 Model Configuration
B.2 Time Consumption
During the dataset construction, query generation utilized 16 H800 GPUs and took 24 hours to process the 30 billion document tokens in The Pile dataset. Following this, 128 CPUs were employed for distributed processing: it took 0.5 hours to generate the keywords, and 3 hours to construct the inverted index based on these keywords. Additionally, since the process is highly parallelizable, scaling it up would not pose significant challenges.
@@ B.3 Filtering of Keywords
B.4 Impact of Split Ratio
This section examines the impact of split_ratio, which controls the proportion of oversampled keyword-based inverted indexes. Figure 9 shows that performance initially improves but then declines as split_ratio increases. The best results occur when oversampled indexes comprise 10-30%. This suggests that moderate oversampling of indexes with fewer documents benefits long-context modeling, highlighting the importance of balanced data distribution for long-context tasks.
@@ B.4 Impact of Split Ratio
Figure 12: Visualizing s Comprising a 128k Context.